OpenStudy (anonymous):

Subtract, and then simplify, if possible. b − 3/b + 9 − b + 2/b − 9

OpenStudy (haleyelizabeth2017):

multiply the first whole fraction by b-9 and the whole second fraction by b+9, I'll show you what you should get :)

OpenStudy (anonymous):

okay

OpenStudy (haleyelizabeth2017):

\[\frac{(b-3)(b-9)}{(b+9)(b-9)}-\frac{(b+9)(b+2)}{(b+9)(b-9)}\] Okay? We needed to get a common denominator, which is \[(b+9)(b-9)\]

OpenStudy (haleyelizabeth2017):

We need to expand the numerators in order to subtract them.

OpenStudy (anonymous):

for the first one I got b^2

OpenStudy (anonymous):

I got b^2 -12b +27 as the numerator

OpenStudy (haleyelizabeth2017):

the first one is \[b^2-12b+27\] you are correct

OpenStudy (haleyelizabeth2017):

And for the second?

OpenStudy (anonymous):

and the second one I got b^2 +11b+18

OpenStudy (haleyelizabeth2017):

right. Now, we have one fraction.\[\frac{(b^2-12b+27)-(b^2+11b+18)}{(b+9)(b-9)}\]

OpenStudy (haleyelizabeth2017):

So, we then have \[\frac{-23b+9}{(b+9)(b-9)}\]

OpenStudy (anonymous):

that's what I got

OpenStudy (haleyelizabeth2017):

Awesome. I don't think we can simplify, but let me check.

OpenStudy (anonymous):

Thankyou

OpenStudy (haleyelizabeth2017):

You're welcome! That's as far as we can simplify haha